# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ A.3 Recursive computation of attributes in the annotation graph
Using the knowledge graph supporting the stoichiometric model of iCH360 (Section 3 in the man text), a number of useful properties can be computed based on simple recursive operations. Here, we outline how such an approach was used to compute i) the molecular mass of all enzyme complexes in the model, based on known masses for all polypeptides and ii) the computation of boolean rules (GPRs) linking all reaction and protein nodes in the graph to the model genes. For a description of the different types of nodes and edges mentioned in this section, see Supplementary Tables 1 and 2.
@@ A.3.1 Computation of molecular masses for all protein nodes
In order to compute the molecular mass of all protein nodes in the graph, the protein nodes corresponding to polypeptides were first annotated with their molecular masses, readily available on the EcoCyc database. Then the molecular mass of all other protein nodes is estimated recursively as follows. Let $\mathcal{I}$ denotes the index set of all protein nodes and $\mathcal{C}(i)\in\mathcal{I}$ the index set of protein components of node $i$, i.e. all nodes connected to node $i$ by a subunit composition relationship. The molecular mass of any protein node $i$, $M_{i}$, is computed as:
@@ A.3.2 Computation of gene-protein-reaction rules
Similarly, boolean gene-protein-reaction (GPR) rules can also be computed recursively. We compute the GPR expression of a node depending on its type (reaction, protein, or logical) and the type of edges connecting it to its neighbouring nodes. Particularly, the GPR expression of node $i$, $G_{i}$, is computed according to the following rules:
  • If node $i$ is a logical AND (logical OR) node, its GPR is the AND (OR) of the GPR of its children nodes.
  • If node $i$ is a polypeptide, its GPR expression is, trivially, the gene associated to the node.
  • If node $i$ is a multimeric protein, its GPR is computed as the AND of the GPR expressions of the children nodes connected to $i$ by a ”subunit composition” edge.
  • If node $i$ is a modified protein, its GPR is computed as the AND of the children nodes connected to $i$ by a ”protein modification” or ”protein modification requirement” edge.
  • If node $i$ is a reaction, its GPR is first constructed as the OR of the GPRs of its catalytic children (nodes connected to $i$ via a catalytic edge, representing isoenzymes for the reaction). To account for non-catalytic protein requirements of the reaction this expression is then ANDed with the GPR expressions of all non-catalytic children (nodes connected to $i$ via a ”non-catalytic requirement” edge). More formally, denoting with $\mathcal{E}(i)$ and $\mathcal{N}(i)$ the index sets of catalytic and non-catalytic children nodes of $i$, respectively, we have:
    $G_{i}=\underbrace{\left[~{}\bigvee_{j\in\mathcal{E}(i)}G_{j}~{}\right]}_{\text{catalytic}}\wedge\underbrace{\left[~{}\bigwedge_{k\in\mathcal{N}(i)}G_{k}~{}\right]}_{\text{non-catalytic}}$ (18)
@@ A.4 Estimation of enzyme complex abundances
In order to estimate the abundance of all enzyme in the model from proteomics data, we use the model graph (Main Text, Section 4.3) to construct a stoichiometric map between enzymes and polypeptides. this map takes the form of a matrix ${\rm\mathbf{E}}\in\mathbb{R}^{n\times m}$, where $n$ is the number of enzymes in the model and $m\geq n$ the number of polypeptides, such that ${\rm\mathbf{E}}_{i,j}$ denotes the stoichiometry of polypeptide $j$ in enzyme $i$. Some polypeptides may be part of additional complexes which are not part of the model. Hence, using the available annotation to the EcoCyc database, we identify these instances and construct a matrix ${\rm\mathbf{\hat{E}}}$ by augmenting ${\rm\mathbf{E}}$ with additional rows corresponding to these out-of-model complexes.
With this mapping at hand, we assume that polypeptide abundances, $\mathbf{p}$ are related to enzyme abundances $\mathbf{e}$ (including the required additional complexes not accounted for in the model) by
@@ A.5 Adjustment of turnover numbers based on proteomics measurements across conditions
In this section, we outline the procedure used to adjust the turnover numbers in EC-iCH360 by fitting proteomic measurements across conditions (see Section 2.5 in the main text). Briefly, our aim is to adjust the turnover numbers that parametrise the model so that enzyme allocation predictions obtained through the enzyme-constrained formulation of FBA (see Methods in the main text) more closely match experimental measurements of enzyme abundances. By simultaneously fitting experimental measurements across many growth conditions, we improve the robustness of the fitting procedure to experimental error and generate a condition-independent set of ”typical” apparent turnover numbers that predict average trends of enzyme allocation across conditions. The output of our procedure is a set of typical enzyme efficiencies, one for each enzyme, estimated from proteomic data across conditions, as well as a set of condition-specific scaling factors that account for differences in total measured enzyme abundances across conditions. In section A.5.1 we rigorously define these parameters and state the main assumption underlying our heuristic. In sections A.5.3, A.5.4, and A.5.5 we formulate a two-steps optimisation problem whose solution, upon a suitable reparameterisation, yields data-fitted estimates of the desired parameters.
@@ A.5.1 Preliminaries
Consider enzyme $i$ in a given metabolic state $j$. This enzyme catalyses a metabolic flux $v_{ij}$ given by
$v_{ij}=\kappa_{ij}~{}c^{enz}_{ij}$ (21)
where $c^{enz}_{ij}$ is the enzyme concentration and $\kappa_{ij}$ is the enzyme efficiency (or “apparent turnover number”). The efficiency $\kappa_{ij}$ is a positive rate (here in 1/h, but usually reported in 1/s). Since, by definition, it must be lower than the enzyme’s turnover number $k_{cat,i}$, we write it as
$\kappa_{ij}=\sigma_{ij}~{}k_{cat,i}$ (22)
where $\sigma_{ij}\in[0,1]$ is a unitless “capacity usage” factor. In enzyme-constrained models, enzymes are often expressed by their mass abundance (in g/gDW) instead of enzyme concentrations, so we can write the flux as
@@
where $a_{ij}$ is the enzyme cost per catalysed flux, given by the enzyme molecular mass $M_{i}$ (in g/mol) divided by the enzyme efficiency $\kappa_{ij}$.
In principle, the capacity usage factors (and therefore efficiencies) of enzymes may freely vary between growth conditions. However, as a heuristics, we here assume that $\sigma_{ij}$ can be approximated as the product of two factors: an enzyme-specific term and a condition-specific one, that is:
@@
$\kappa_{ij}=\underbrace{\sigma_{i}~{}k_{cat,i}}_{\kappa_{i}}~{}\tau_{j}$ (25)
where $\kappa_{i}$ is a ”typical apparent turnover number” that is condition-independent. Practically, the above assumption allows us to simplify the problem by reducing the number of parameters to be fitted from $I\cdot J$ to $I+J$, where $I$ and $J$ denote the total number of enzymes and conditions considered, respectively. Our heuristic assumption corresponds to the idea that ”high-quality carbon sources” allow a cell to establish metabolic states in which enzyme efficiencies are generally high, allowing for large fluxes per enzyme abundance in all the reactions, and therefore for high cell growth rates. Probably, this heuristics would fail in some other cases, e.g. cases in which enzymes are specifically perturbed by enzyme inhibitors. But in fact, it turns out that our model, assuming a single ”typical apparent turnover number” $k_{app,i}$ for each enzyme, yields very good enzyme allocation predictions. This is exactly what would be expected if our heuristic assumption were correct, and therefore supports our heuristic prediction.
Below we describe how the estimates of (enzyme-specific) efficiencies $\kappa_{i}$ and (condition specific) scaling factors $\tau_{j}$ for our model were obtained from model simulations and proteomics data.
@@ A.5.2 Overview of the fitting procedure
Fitting the typical turnover parameters to proteomic data, in general, not simple. Due to the linear programming formulation underlying the enzyme-constrained FBA problem, optimal flux distributions (and, by direct consequence, enzyme allocation predictions) are discontinuous over the turnover parameter space, making derivative-based searches through this space problematic from a numerical perspective. Similarly, the high dimensionality of the parameter space limits the applicability of gradient-free optimisation algorithms. Hence, we restrict ourselves to the (comparatively simpler) problem of adjusting turnover parameters for a fixed set of reference flux distributions across growth conditions, constraining our search to the portion of parameter space in which these reference flux vectors remain optimal for their respective conditions. This simplified fitting procedure thus consists of two steps. In the first part of the procedure, we use an initial parameter set to compute a set of reference flux distributions (one per growth condition) using enzyme-constrained FBA. In the second part, turnover parameters are fitted based on these reference flux distributions and experimental measurements of enzyme abundances.
@@ A.5.3 Obtaining a set of reference flux distributions
Following the sMOMENT formulation of enzyme-constrained FBA [bekiaris_automatic_2020], we consider a metabolic network with $N$ reactions and $M$ metabolites where all metabolic fluxes are positive (i.e. reversible reactions are split into forward and backwards components) and at most one enzyme is associated with each reaction. Hence, we assume that the enzyme cost required to sustain flux $v_{i}$ for a given growth condition $j$ is given by $a_{ij}~{}v_{i}$, where the cost per unit flux $a_{ij}$ is given by:
@@
Here, $M_{i}$ is the molecular mass of the enzyme associated with the reaction, $\kappa_{ij}$ is the condition-dependent enzyme efficiency, as defined in 25. Given an initial guess for the value of each $\kappa_{ij}$, we compute a reference flux distribution for the $j$th growth condition, $\mathbf{v}_{j}^{*}$ by fixing the biomass flux, $v_{\rm BM}$ to the experimentally measured rate and minimizing the total enzyme cost:
@@
Here, the objective $\mathbf{a}_{j}^{\top}\mathbf{v}$ is the total enzyme cost for this condition, $\bar{v}_{\mathrm{BM},j}$ is the experimentally measured growth rate for the $j$th growth condition, ${\rm\mathbf{S}}\in\mathbb{R}^{\rm{M}\times\rm{N}}$ is the stoichiometric matrix of the network and ${\rm\mathbf{B}}_{j}\in\mathbb{R}^{\rm{P}\times\rm{N}}$ and $\mathbf{b}_{j}\in\mathbb{R}^{\rm{P}}$ are a matrix and a vector, respectively, encoding any desired upper bound (or positive lower bound) on the fluxes for the $j$th growth condition. Noting that constraint 27c can be equivalently cast as a double inequality, we rewrite the problem in the more general form:
@@
$\rm{s.t.}$ $\quad{\rm\mathbf{S}}\mathbf{v}=\mathbf{0}$ $\qquad\qquad\rm{(a)}$
$\quad\hat{{\rm\mathbf{B}}}_{j}\mathbf{v}\leq\hat{\mathbf{b}}_{j}$ $\qquad\qquad\rm{(b)}$
@@
In order to solve the the linear program 27, we shall consider an initial guess for the turnover parameters and assume that, for each condition, all enzymes operate at the same saturation level, so that $\sigma_{ij}\equiv\bar{\sigma}_{j}$. Since the optimal flux distribution obtained as a solution of problem (28) is unchanged by the choice of $\bar{\sigma}_{j}$ (as this merely amounts to a scaling of the objective function), we can simply set $\bar{\sigma}_{j}=1$ (that is, set $\kappa_{ij}=k_{\rm cat,i}$) for all conditions at this stage.
@@ A.5.4 Estimating typical enzyme efficiencies for the reference set of flux distributions
We now turn to fitting the relevant parameters against experimental measurements of enzyme abundances. For this purpose, we shall express the the efficiency of each enzyme-condition pair as:
@@
where $\theta_{j}$ is a condition-specific scaling term that simultaneously scales all efficiencies for a given condition, while $k_{i}$ is an ”adjusted” turnover parameter that simultaneously accounts for inaccuracies in the original turnover parameter numbers as well as differences in typical saturation across enzymes (and, hence, it’s not formally a turnover number). While the above parameterization differs from the one in equation 25 – the terms have a different interpretation! – it will greatly simplify the notation of the fitting problem, as it allows us to easily distinguish between ”global” adjustments (those affecting all enzymes in a condition), which we wish to pick freely, and ”local” adjustments (affecting the efficiency of individual enzymes), which instead we wish to regularise. While this factorisation of saturation effects is merely a computational convenience and not necessarily biologically meaningful, we will retrieve the parameters in Eq. 25 from those in Eq. 30 upon a simple reparameterisation, as we detail in Section A.5.5
@@
$\rm{s.t.}$ $\quad\mathbf{l}\leq\mathbf{p}-\bar{\mathbf{p}}\leq\mathbf{u}$
@@
where $N_{\rm e}$ is the number of enzyme-condition pairs, $N_{\rm p}$ is the number of turnover parameters, $\bar{e}_{ij}$ is the experimental measurement of enzyme $i$ in condition $j$, $\mathbf{l}$ and $\mathbf{u}$ are bounds on the allowable adjustment, $\rho>0$ is a scalar hyperparameter, and the function $l(\cdot)$ is defined as:
@@
In order to define the region $\mathcal{S}$, we shall exploit the sufficient optimality conditions of LP (28). Introducing, for each growth condition, the vectors of dual variables, $\boldsymbol{\lambda}_{j}\in\mathbb{R}^{M}$ and $\boldsymbol{\mu}_{j}\in\mathbb{R}^{P}$, corresponding to constraints 28b and 28c, respectively, we note that problem 28 admits (for the $j$th growth condition) a dual problem in the form:
$\max_{\boldsymbol{\lambda}_{j},~{}\boldsymbol{\mu}_{j}}$ $\quad-\mathbf{b}_{j}^{\top}\boldsymbol{\mu}_{j}$ (34)
$\rm{s.t.}$ $\quad{\rm\mathbf{S}}^{\top}\boldsymbol{\lambda}_{j}~{}+{\rm\mathbf{B}}_{j}^{\top}\boldsymbol{\mu}_{j}+\mathbf{a}\geq\mathbf{0}$ $\qquad\qquad\rm{(a)}$
$\quad\boldsymbol{\mu}_{j}\geq\mathbf{0}$
A well-known result in linear programming duality theory [maranas_zomorrodi_2016] states that a flux distribution is the optimal solution of the primal problem 28, if, and only if, the dual problem 34 is feasible (dual feasibility) and its optimal objective coincides with the primal optimal objective, that is $\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\mathbf{b}_{j}^{\top}\boldsymbol{\mu}_{j}$ (strong duality). Taken together, dual feasibility and strong duality thus define the region of optimality in turnover parameter space of each reference flux distribution. Hence, we can integrate the above definition of $\mathcal{S}$ within problem 32 by introducing the two vectors of dual variables ($\boldsymbol{\lambda}_{j}$ and $\boldsymbol{\mu}_{j}$) for each condition as additional optimisation variables, and simultaneously enforcing the optimality for each reference distribution. By doing this, we obtain the final formulation of the nonlinear program for turnover number adjustment, which we solved to obtain the results shown in the main text:
$\rm{s.t.}$ $\quad\mathbf{l}\leq\mathbf{p}-\bar{\mathbf{p}}\leq\mathbf{u}$
$\quad{\rm\mathbf{S}}^{\top}\boldsymbol{\lambda}_{j}~{}+{\rm\mathbf{B}}_{j}^{\top}\boldsymbol{\mu}_{j}+\mathbf{a}_{j}\geq\mathbf{0}$ $\quad j=1,\dotsc,J$
$\quad\boldsymbol{\mu}_{j}\geq\mathbf{0}$ $\quad j=1,\dotsc,J$
$\quad\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\mathbf{b}_{j}^{\top}\boldsymbol{\mu}_{j}$ $\quad j=1,\dotsc,J$

### v3

@@ A.3 Computation of attributes using the knowledge graph
Using the knowledge graph supporting the stoichiometric model of iCH360 (Section 3 in the man text), a number of useful properties can be computed based on simple operations. Here, we outline how such an approach was used to i) compute the molecular mass of all enzyme complexes in the model, based on known masses for all polypeptides and ii) construct the boolean rules (GPRs) linking all reactions and proteins in the graph to the model genes. For a description of the different types of nodes and edges mentioned in this section, see Supplementary Tables 2 and 3.
@@ A.3.1 Computation of molecular masses for all protein nodes
In order to compute the molecular masses of all protein nodes in the graph, the protein nodes corresponding to polypeptides were first annotated with their molecular masses, readily available on the EcoCyc database. Then the molecular masses of all other protein nodes are estimated recursively as follows. Let $\mathcal{I}$ denotes the index set of all protein nodes and $\mathcal{C}(i)\in\mathcal{I}$ the index set of protein components of node $i$, i.e. all nodes connected to node $i$ by a subunit composition relationship. The molecular mass of any protein node $i$, $M_{i}$, is computed as:
@@ A.3.2 Computation of gene-protein-reaction rules
Boolean gene-protein-reaction (GPR) rules are a widely used tool defining a map between a genotype (set of active genes) and a phenotype (set of active reactions) in a metabolic model. Conventionally, this is achieved by assigning to each reaction a boolean expression given in terms of genes in the model. In this section, we show how such expressions were computed for the reactions in iCH360 using the knowledge graph.
Starting from the leaves of the graph (genes), we construct, for each node, a boolean expression describing its state (active/inactive, corresponding to a boolean true / false) in terms of its children. The exact form of this boolean expression depends on the type of the node (reaction, protein, or logical) and the type of edges connecting it to its neighbours (Figure S22). Particularly:
  • A polypeptide node is active if its associated gene is also active (Figure S22A, left).
  • A multimeric protein is active if all of its subunits (the child nodes connected to it by a ”subunit composition” edge) are active (Figure S22A, middle).
  • A modified protein is active if its unmodified form (the child node connected to it by a ”protein modification” edge) and its modification requirements (the child nodes connected to it by a ”protein modification requirement” edge) are active (Figure S22A, right).
  • A logical AND (logical OR) node is active if all (any) of its child nodes are active (Figure S22B).
  • Finally, a reaction node is active if any of its catalysing isoenzymes (the child nodes connected to it via a ”catalysis” edge) are active and, at the same time, all of its catalytic requirements (the child nodes connected to it via a ”non-catalytic requirement” edge) are also active (Figure S22C).
Using these definitions, the boolean expression describing the state of a reaction can be written, ultimately, solely in terms of genes, enabling computation of conventional GPR rules and their incorporation in the standard metabolic model.
@@ A.4 Estimation of enzyme complex abundances
In order to estimate the abundances of all enzymes in the model from proteomics data, we use the model graph (Main Text, Section 4.4) to construct a stoichiometric map between enzymes and polypeptides. This map takes the form of a matrix ${\rm\mathbf{E}}\in\mathbb{R}^{n\times m}$, where $n$ is the number of enzymes in the model and $m\geq n$ the number of polypeptides, such that ${\rm\mathbf{E}}_{ij}$ denotes the stoichiometry of polypeptide $j$ in enzyme $i$. Some polypeptides may be part of additional enzyme complexes that are not part of the model. Using the available annotation to the EcoCyc database, we identified $7$ such polypeptides, mapping to $9$ out-of-model complexes. If these out-of-model complexes were not taken into account, the abundance of model enzymes to which these polypeptides map would be overestimated. Hence, we constructed a matrix ${\rm\mathbf{\hat{E}}}$ by augmenting ${\rm\mathbf{E}}$ with additional rows corresponding to the identified out-of-model complexes.
With this mapping at hand, we assume that polypeptide abundances $\mathbf{p}$ are related to enzyme abundances $\mathbf{e}$ (including the required additional complexes not accounted for in the model) by
@@ A.5 Adjustment of turnover numbers based on proteomics measurements across conditions
In this section, we outline the procedure used to adjust the turnover numbers in EC-iCH360 by fitting proteomic measurements across conditions (see Section 2.4 in the main text). Briefly, our aim is to adjust the turnover numbers that parametrise the model so that enzyme allocation predictions obtained through the enzyme-constrained formulation of FBA (see Methods in the main text) match more closely experimental measurements of enzyme abundances. By simultaneously fitting experimental measurements across many growth conditions, we improve the robustness of the fitting procedure to experimental error and generate a condition-independent set of ”typical” apparent turnover numbers that predict average trends of enzyme allocation across conditions. The output of our procedure is a set of typical enzyme efficiencies, one for each enzyme, estimated from proteomic data across conditions, as well as a set of condition-specific scaling factors that account for differences in total measured enzyme abundances between conditions. In section A.5.1 we rigorously define these parameters and state the main assumption underlying our heuristic. In sections A.5.3, A.5.4, and A.5.5 we formulate a two-steps optimisation problem whose solution, upon a suitable reparameterisation, yields data-fitted estimates of the desired parameters.
@@ A.5.1 Preliminaries
Consider an enzyme $i$ in a given metabolic state $j$. This enzyme catalyses a metabolic flux $v_{ij}$ (in (mol/gDW)/h) given by
$v_{ij}=\kappa_{ij}~{}c^{\rm enz}_{ij}$ (20)
where $c^{\rm enz}_{ij}$ is the enzyme concentration and $\kappa_{ij}$ is the enzyme efficiency (or “apparent turnover number $k_{\rm app}$”). The efficiency $\kappa_{ij}$ is a positive rate (here in 1/h, but usually reported in 1/s). Since, by definition, it must be lower than the enzyme’s turnover number $k_{\rm{cat},i}$, we write it as
$\kappa_{ij}=\sigma_{ij}~{}k_{\rm{cat},i}$ (21)
where $\sigma_{ij}\in[0,1]$ is a unitless “capacity usage” factor. In enzyme-constrained models, enzymes are often expressed by their mass abundance $e_{i}=M_{i}~{}c^{\rm enz}_{i}$ (in g/gDW) instead of enzyme concentrations, where $M_{i}$ is the enzyme molecular mass (in g/mol), so we can write the flux as
@@
where $a_{ij}$ is the enzyme cost per catalysed flux, given by the molecular mass $M_{i}$ divided by the enzyme efficiency $\kappa_{ij}$.
In principle, the capacity usage factors $\sigma_{ij}$ (and therefore efficiencies) of enzymes may freely vary between growth conditions. However, as a heuristic, we here assume that they can be approximated as the product of two factors: an enzyme-specific term and a condition-specific one, that is:
@@
$\kappa_{ij}=\underbrace{\sigma_{i}~{}k_{\rm{cat},i}}_{\kappa_{i}}~{}\tau_{j}$ (24)
where $\kappa_{i}$ is a ”typical apparent turnover number” that is condition-independent. Practically, the above assumption allows us to simplify the problem by reducing the number of parameters to be fitted from $I\cdot J$ to $I+J$, where $I$ and $J$ denote the total number of enzymes and conditions considered, respectively. Our heuristic assumption corresponds to the idea that ”high-quality carbon sources” allow a cell to establish metabolic states in which enzyme efficiencies are generally high, allowing for large fluxes per enzyme abundance in all the reactions, and therefore for high cell growth rates. Probably, this heuristics would fail in some other cases, e.g. cases in which enzymes are specifically perturbed by enzyme inhibitors. But in fact, it turns out that our model, assuming a single ”typical apparent turnover number” $k_{\rm{app},i}$ for each enzyme, yields very good enzyme allocation predictions. This is what would be expected if our heuristic assumption were correct, and therefore supports our heuristic prediction.
We now describe how the estimates of (enzyme-specific) efficiencies $\kappa_{i}$ and (condition specific) scaling factors $\tau_{j}$ for our model were obtained from model simulations and proteomics data.
@@ A.5.2 Overview of the fitting procedure
Fitting the typical turnover parameters to proteomic data is, in general, not simple. Due to the linear programming formulation underlying the enzyme-constrained FBA problem, optimal flux distributions (and, by direct consequence, enzyme allocation predictions) are discontinuous over the turnover parameter space, making derivative-based searches through this space problematic from a numerical perspective. Similarly, the high dimensionality of the parameter space limits the applicability of gradient-free optimisation algorithms. Hence, we restrict ourselves to the (comparatively simpler) problem of adjusting turnover parameters for a fixed set of reference flux distributions across growth conditions, constraining our search to the portion of parameter space in which these reference flux vectors remain optimal for their respective conditions. This simplified fitting procedure thus consists of two steps. In the first part of the procedure, we use an initial parameter set to compute a set of reference flux distributions (one per growth condition) using enzyme-constrained FBA. In the second part, turnover parameters are fitted based on these reference flux distributions and experimental measurements of enzyme abundances.
@@ A.5.3 Obtaining a set of reference flux distributions
Following the sMOMENT formulation of enzyme-constrained FBA [bekiaris_automatic_2020], we consider a metabolic network with $N$ reactions and $M$ metabolites where all metabolic fluxes are positive (i.e. reversible reactions are split into forward and backwards components) and at most one enzyme is associated with each reaction (see Methods in the main text for more information about how such unique mapping between reactions and catalyzing enzymes was generated in our case). Hence, we assume that the enzyme cost required to sustain flux $v_{i}$ for a given growth condition $j$ is given by $a_{ij}~{}v_{i}$, where the cost per unit flux $a_{ij}$ is given by:
@@
Here, $M_{i}$ is the molecular mass of the enzyme associated with the reaction, $\kappa_{ij}$ is the condition-dependent enzyme efficiency, as defined in 24. Given an initial guess for the value of each $\kappa_{ij}$, we compute a reference flux distribution for the $j$th growth condition, $\mathbf{v}_{j}^{*}$ by fixing the biomass flux, $v_{\rm BM}$ to the experimentally measured rate and minimizing the total enzyme cost:
@@
Here, the objective $\mathbf{a}_{j}^{\top}\mathbf{v}$ is the total enzyme cost for the $j$th growth condition, $\bar{v}_{\mathrm{BM},j}$ is the experimentally measured growth rate for the condition, ${\rm\mathbf{S}}\in\mathbb{R}^{\rm{M}\times\rm{N}}$ is the stoichiometric matrix of the network and ${\rm\mathbf{B}}_{j}\in\mathbb{R}^{\rm{P}\times\rm{N}}$ and $\mathbf{b}_{j}\in\mathbb{R}^{\rm{P}}$ are a matrix and a vector, respectively, encoding any desired upper bound (or positive lower bound) on the fluxes for the growth condition. Noting that constraint 26c can be equivalently cast as a double inequality, we rewrite the problem in the more general form:
@@
$\rm{s.t.}$ $\quad{\rm\mathbf{S}}~{}\mathbf{v}=\mathbf{0}$ $\qquad\qquad\rm{(a)}$
$\quad\hat{{\rm\mathbf{B}}}_{j}~{}\mathbf{v}\leq\hat{\mathbf{b}}_{j}$ $\qquad\qquad\rm{(b)}$
@@
In order to solve the the linear program 26, we shall consider an initial guess for the turnover parameters and assume that, for each condition, all enzymes operate at the same saturation level, so that $\sigma_{ij}\equiv\bar{\sigma}_{j}$. Since the optimal flux distribution obtained as a solution of problem (27) is unchanged by the choice of $\bar{\sigma}_{j}$ (as this merely amounts to a scaling of the objective function), we can simply set $\bar{\sigma}_{j}=1$ (that is, set $\kappa_{ij}=k_{\rm cat,i}$) for all conditions at this stage.
@@ A.5.4 Estimating typical enzyme efficiencies for the reference set of flux distributions
We now turn to fitting the relevant parameters against experimental measurements of enzyme abundances. For this purpose, we shall express the efficiency of each enzyme-condition pair as:
@@
where $\theta_{j}$ is a condition-specific scaling term that simultaneously scales all efficiencies for a given condition, while $k_{i}$ is an ”adjusted” turnover parameter that simultaneously accounts for inaccuracies in the original turnover parameter numbers as well as differences in typical saturation across enzymes (and, hence, it’s not formally a turnover number). While the above parameterization differs from the one in equation (24) – the terms have a different interpretation! – it will greatly simplify the notation of the fitting problem, as it allows us to easily distinguish between ”global” adjustments (those affecting all enzymes in a condition), which we wish to pick freely, and ”local” adjustments (affecting the efficiency of individual enzymes), which instead we wish to regularise. While this factorisation of saturation effects is merely a computational convenience and not necessarily biologically meaningful, we will retrieve the parameters in Eq. (24) from those in Eq. (29) upon a simple reparameterisation, as we detail in Section A.5.5
@@
$\rm{s.t.}$ $\quad\mathbf{u_{\rm min}}\leq\mathbf{p}-\bar{\mathbf{p}}\leq\mathbf{u_{\rm max}}$
@@
where $N_{\rm e}$ is the number of enzyme-condition pairs, $N_{\rm p}$ is the number of turnover parameters, $\bar{e}_{ij}$ is the experimental measurement of enzyme $i$ in condition $j$, $\mathbf{u_{\rm min}}$ and $\mathbf{u_{\rm max}}$ are bounds on the allowable adjustment, $\rho>0$ is a scalar hyperparameter, and the function $l(\cdot)$ is defined as:
@@
In order to define the region $\mathcal{S}$, we shall exploit the sufficient optimality conditions of LP (27). Introducing, for each growth condition, the vectors of dual variables, $\bm{\lambda}_{j}\in\mathbb{R}^{M}$ and $\bm{\mu}_{j}\in\mathbb{R}^{P}$, corresponding to constraints 27b and 27c, respectively, we note that problem (27) admits (for the $j$th growth condition) a dual problem in the form:
$\max_{\bm{\lambda}_{j},~{}\bm{\mu}_{j}}$ $\quad-\hat{\mathbf{b}}_{j}^{\top}\bm{\mu}_{j}$ (33)
$\rm{s.t.}$ $\quad{\rm\mathbf{S}}^{\top}\bm{\lambda}_{j}~{}+\hat{{\rm\mathbf{B}}}_{j}^{\top}\bm{\mu}_{j}+\mathbf{a}_{j}\geq\mathbf{0}$ $\qquad\qquad\rm{(a)}$
$\quad\bm{\mu}_{j}\geq\mathbf{0}$
A well-known result in linear programming duality theory [maranas_zomorrodi_2016] states that a flux distribution is the optimal solution of the primal problem (27), if, and only if, the dual problem (33) is feasible (dual feasibility) and its optimal objective coincides with the primal optimal objective, that is $\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\mathbf{b}_{j}^{\top}\bm{\mu}_{j}$ (strong duality). Taken together, dual feasibility and strong duality thus define the region of optimality in turnover parameter space of each reference flux distribution. Hence, we can integrate the above definition of $\mathcal{S}$ within problem (31) by introducing the two vectors of dual variables ($\bm{\lambda}_{j}$ and $\bm{\mu}_{j}$) for each condition as additional optimisation variables, and simultaneously enforcing the optimality for each reference distribution. By doing this, we obtain the final formulation of the nonlinear program for turnover number adjustment, which we solved to obtain the results shown in the main text:
$(\mathbf{p}^{*},\mathbf{s}^{*})=\operatorname*{arg\,min}_{\mathbf{p},~{}\mathbf{s}}$ $\quad\frac{1}{N_{\rm e}}\sum_{ij}~{}\big{[}~{}l(e_{ij})-l(\bar{e}_{ij})~{}\big{]}^{2}+\frac{\rho}{N_{\rm p}}\sum_{i}\big{(}~{}p_{i}-\bar{p}_{i}~{}\big{)}^{2}$ (34)
$\rm{s.t.}$ $\quad\mathbf{u_{\rm min}}\leq\mathbf{p}-\bar{\mathbf{p}}\leq\mathbf{u_{\rm max}}$
$\quad{\rm\mathbf{S}}^{\top}\bm{\lambda}_{j}~{}+\hat{{\rm\mathbf{B}}}_{j}^{\top}\bm{\mu}_{j}+\mathbf{a}_{j}\geq\mathbf{0}$ $\quad j=1,\dotsc,J$
$\quad\bm{\mu}_{j}\geq\mathbf{0}$ $\quad j=1,\dotsc,J$
$\quad\mathbf{a}_{j}^{\top}\mathbf{v}_{j}^{*}=-\hat{\mathbf{b}}_{j}^{\top}\bm{\mu}_{j}$ $\quad j=1,\dotsc,J$